For Firm A, when four units of output are produced, the total cost is $175 and the average variable cost is $33.75. What would the average fixed cost be if the units were produced?
Answer #1

Solution -

Total cost = 175

Total variable cost 4 units

= 33.75*4

=135

Now Calculate

TFC = TC - TVC

= 175 -135

=40

Now Calculate

Average fixed cost of = 40/10

= $4